OpenStudy (anonymous):

The velocity of a particle moving in a straight line is given by the following formula. V=3te^(t^2)+3t (a) Find an expression for the position s after a time t. s=? (b) Given that s = 6 at time t = 0, find the constant of integration C and hence an expression for s in terms of t without any unknown constants. s=

OpenStudy (valpey):

This?:\[V=3te^{t^2}+3t\]

OpenStudy (valpey):

\[\int{due^u=e^u+C};\ \ Let \ u=t^2\]

OpenStudy (anonymous):

yes then what would the answers be?

OpenStudy (anonymous):

You tell us what it is and we can check if it's right or wrong.

OpenStudy (anonymous):

me? i have no clue what it is